The current asset section of the Excalibur Tire Company’s balance sheet consists of cash, marketable securities, accounts receivable, and inventory. The December 31, 2021, balance sheet revealed the following: Inventory $ 990,000 Total assets $ 3,500,000 Current ratio 2.10 Acid-test ratio 1.20 Debt to equity ratio 1.8 Required: Determine the following 2021 balance sheet items:

1) Current Assets

2) Shareholder Equity

3) Long-term assets

4) Long-term liabilities
Answer #1

1. Quick assets = current asset- inventories = current asset - 990,000

Quick Ratio = Quick assets / current liabilities = 1.20

Current Ratio = Current assets / current liabilities = 2.10

Quick Ratio = current asset - 990,000 / current liabilities = 1.20

990,000 / current liabilities = 0.90

current liabilities = $1100000

Current assets = 1100000*2.10 = $2310000

2. Debt to equity ratio = total liabilities / shareholders equity = 1.80

Total liabilities + shareholders equity = Total assets

Total liabilities + shareholders equity = $3,500,000

Let x will be the shareholders equity

1.80x + x = 3500000

shareholders equity = $1250000

3. Long-term assets = Total asset - current asset

= $3,500,000 - $2310000 = $1190000

4. Long term liabilities = total assets- current liabilities - share holders equity

= 3500000 - 1100000 - 1250000

= $1150000